A single card is drawn from a well-shuffled deck. This experiment is performed 5 times. (The card is replaced and the deck is shuffled between repetitions of the experiment.) What is the probability that exactly 2 spades and exactly one diamond are drawn?

A.

B.

C.

D.

E.
Answer #1

probability that exactly 2 spades and exactly one diamond are drawn

=(5!/(2!*1!*2!)) *(1/4)2*(1/4)1*(1/2)2 =15/128 =0.1172